# Robust Single Rotation Averaging Revisited

### v4

@@ IV METHOD
$\mathbf{R}_{\mathrm{init}}\leftarrow\operatorname*{\arg\!\min}_{\displaystyle\mathbf{R}_{j\in\{1\cdots N\}}}\sum_{i=1}^{N}\min\left(\epsilon_{\mathrm{c}},\lVert\mathbf{R}_{i}-\mathbf{R}_{j}\rVert_{F}\right)$; $\mathcal{I}\leftarrow\{i\ |\ \lVert\mathbf{R}_{i}-\mathbf{R}_{\mathrm{init}}\rVert_{F}<\epsilon_{\mathrm{c}},i=1,\cdots,N\}$; /* Set the chordal $L_{2}$-mean [14] of inliers as the initial estimate: */
$\mathbf{R}_{\mathrm{avg}}\leftarrow\mathrm{proj}_{SO(3)}\left(\sum_{i\in\mathcal{I}}\mathbf{R}_{i}\right)$; /* Run the Weiszfeld algorithm on SO(3) [4] over the inlier set: */
1 for $\mathrm{it}=1,2,3,\cdots,\mathrm{it}_{\mathrm{max}}$ do
2       $\mathbf{v}_{i}\leftarrow\mathrm{Log}\left(\mathbf{R}_{i}\mathbf{R}_{\mathrm{avg}}^{\top}\right)\ \text{for}\ i\in\mathcal{I}$; $\Delta\mathbf{v}\leftarrow\frac{\sum_{i\in\mathcal{I}}\mathbf{v}_{i}/\lVert\mathbf{v}_{i}\rVert}{\sum_{i\in\mathcal{I}}1/\lVert\mathbf{v}_{i}\rVert}$; $\mathbf{R}_{\mathrm{avg}}\leftarrow\mathrm{Exp}(\Delta\mathbf{v})\mathbf{R}_{\mathrm{avg}}$; if $\lVert\Delta\mathbf{v}\rVert<\delta$ then
@@ V RESULTS
We compare our method against three existing methods: Hartley2011 [4], Lee2020 [22] and ROBIN [26]. All experiments are conducted in Matlab on a laptop with i7-4810MQ CPU (2.8 GHz) and 16 GB RAM.

### v5

@@ IV METHOD
1
[…]
3$\mathcal{I}\leftarrow\{i\ |\ \lVert\mathbf{R}_{i}-\mathbf{R}_{\mathrm{init}}\rVert_{F}<\epsilon_{\mathrm{c}},i=1,\cdots,N\}$;
/* Set the chordal $L_{2}$-mean [14] of inliers as the initial estimate: */
4
[…]
/* Run the Weiszfeld algorithm on SO(3) [4] over the inlier set: */
6
7for $\mathrm{it}=1,2,3,\cdots,\mathrm{it}_{\mathrm{max}}$ do
8       $\mathbf{v}_{i}\leftarrow\mathrm{Log}\left(\mathbf{R}_{i}\mathbf{R}_{\mathrm{avg}}^{\top}\right)\ \text{for}\ i\in\mathcal{I}$;
9      $\Delta\mathbf{v}\leftarrow\frac{\sum_{i\in\mathcal{I}}\mathbf{v}_{i}/\lVert\mathbf{v}_{i}\rVert}{\sum_{i\in\mathcal{I}}1/\lVert\mathbf{v}_{i}\rVert}$;
10      $\mathbf{R}_{\mathrm{avg}}\leftarrow\mathrm{Exp}(\Delta\mathbf{v})\mathbf{R}_{\mathrm{avg}}$;
11      if $\lVert\Delta\mathbf{v}\rVert<\delta$ then
@@ V RESULTS
We compare our method against three existing methods: Hartley2011 [4], Lee2020 [22] and ROBIN [26]. All methods are implemented in Matlab and run on a laptop with i7-4810MQ CPU (2.8 GHz) and 16 GB RAM.